Not only was the Stanley Hotel in Estes Park, Colorado the inspiration for Stephen King's classic The Shining, it is also listed among the most haunted locations in the United States. Guests stay at the Stanley hoping to experience the paranormal in the form of ghostly apparitions. Unfortunately, only 20% of the guests actually have these experiences. If 3 visitors to the Stanley are randomly selected, what is the probability that all 3 see a ghost?
Answer #1

Percentage of guests who actually have these experiences = 20%

Thus Probability = 20/100 = 0.2 (of one person)

Probability that all 3 see a ghost = 0.2 x 0.2 x 0.2 = 0.008